### v2

@@ 5.4 $6/7$-MMS with Copies
By Lemma 15, the bags $B_{1},\ldots,B_{n-n^{\prime}}$ are allocated during the round robin phase. By Lemma 17, there exists at least $n^{\prime}$ many bags $B_{i_{1}},\ldots,B_{i_{n^{\prime}}}$ for $i_{1}<i_{2}<\ldots<i_{n^{\prime}}\leq n-n^{\prime}$ that were allocated to other agents during the round robin phase and $v_{i}(B_{i_{\ell}})>1$ for all $\ell\in[n^{\prime}]$. Note that $i_{n^{\prime}}\geq n^{\prime}$. Since $B_{j}$ have been assigned a good in all but maybe the last round, we have $|B_{j}|\geq|B_{i_{n^{\prime}}}|-1$. Let $B_{j}=\{g_{1},\ldots,g_{k}\}$ and $B_{i_{n^{\prime}}}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k^{\prime}}\}$ for $j=g_{1}<\ldots<g_{k}$ and $i_{n^{\prime}}=g^{\prime}_{1}<\ldots<g^{\prime}_{k^{\prime}}$. For all $\ell\in[k^{\prime}-1]$, since $g_{\ell}$ is allocated before $g^{\prime}_{\ell+1}$, $g_{\ell}<g^{\prime}_{\ell+1}$ and hence $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell+1})$. We have
@@
Theorem 4.
Given an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with $n^{*}$ agents and $\alpha\leq 6/7$,
@@
returns a simple $\alpha$-MMS allocation with at most $\lfloor n^{*}/2\rfloor$ many distinct copies.
Proof.
Let $n$ be the number of agents after applying $\mathtt{S-reduce(\mathcal{I},\alpha)}$. Since $S^{\alpha}$ and $R^{\alpha}_{k}$ are valid reductions for all $k<|M|/n^{*}$, all agents who receive a bag during the $\mathtt{S-reduce(\mathcal{I},\alpha)}$ receive at least $\alpha$ fraction of their MMS value and the MMS value of the remaining agents do not decrease after this process. Without loss of generality, let the MMS value of the remaining agents be $1$ at this point of the algorithm. All the agents who receive a bag during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}$ get at least $\alpha$. Let agent $i$ be an agent who receives bag $B_{j}$ after the round robin phase. Let $g$ be the good that was duplicated and added to $B_{j}$ right before allocating it to agent $i$. $g\leq n^{\prime}$ and thus $v_{i}(g)\geq v_{i}(n^{\prime})$. By Lemma 18, $v_{i}(B_{j}\cup\{g\})\geq\alpha$. Therefore, the final allocation is $\alpha$-MMS.
We now prove that the generated MMS allocation is simple, namely that there are $t\leq\lfloor n^{*}/2\rfloor$ distinct copies, and that the $2t$ instances, composed of these copies and their corresponding original goods, are allocated to $2t$ distinct agents. Let $k=n^{*}-n$ be the number of agents who receive a bag in $\mathtt{S-reduce(\mathcal{I},\alpha)}$. Only when applying $S^{\alpha}$, new duplications are introduced and one duplication per two agents that are removed. So during $\mathtt{S-reduce(\mathcal{I},\alpha)}$, at most $\lfloor k/2\rfloor$ distinct copies are used. Each instance of these good is allocated to a different agent. Furthermore, all the instances of the goods that are allocated during $S^{\alpha}$ along with their owners, are removed from the instance. Therefore, these goods cannot be copied more than once and their owners cannot get any other goods later on. In the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}$, we duplicate items $1,2,\ldots,n^{\prime}$ once and allocate the original and the duplications to $2n^{\prime}$ different agents. By Corollary 6, $n^{\prime}\leq\lfloor n/2\rfloor$. Therefore, in total at most $\lfloor n^{*}/2\rfloor$ distinct copies are used and the allocation is simple. ∎
5.5 $4/5$-MMS with Copies
In this section, we present an algorithm which, given an ordered instance, outputs a simple $\alpha$-MMS allocation with at most $\lfloor n/3\rfloor$ distinct copies for any $\alpha\leq 4/5$.
@@ 5.5 $4/5$-MMS with Copies
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
@@
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
@@
The lemma follows from Lemma 13 and Corollary 4. ∎
For the rest of this (sub)-section, let $\alpha\leq 4/5$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible and $T^{\alpha}$-irreducible instance, and $\mu_{i}(\mathcal{I})=1$. Let $n$ be the number of agents right before the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, =$\mathcal{B}=\{B_{1},\ldots,B_{a-1}\}$ be the set of bags allocated during this phase, and $n^{\prime}=n-a+1$ be the number of remaining agents after this phase.
Lemma 20.
@@
Lemma 21.
Let $i$ be a remaining agent after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ (if any). Then |$|\{B\in\mathcal{B}\mid v_{i}(B)>\alpha\}|\geq 2n/3$ or $n\leq 5$.
@@
Assume $n\geq 6$. Let +=$\mathcal{B}^{+}=\{B\in\mathcal{B}\mid v_{i}(B)>\alpha\}$. By definition, for all $B\in\mathcal{B}\setminus\mathcal{B}^{+}$, $v_{i}(B)\leq\alpha$. Let $\bar{\mathcal{B}}$ be the set of $n-|\mathcal{B}|=n^{\prime}$ many remaining bags after the round robin phase. For all $B\in\bar{\mathcal{B}}$, $v_{i}(B)<\alpha$. Otherwise, $B$ should have been allocated to $i$ (or some other agent who values it at least $\alpha$ fraction of their MMS value). We consider two cases:
@@
For all $B\in\mathcal{B}\setminus\{B_{1},B_{2}\}$, Lemma 20 implies $|B|\geq 3$ and Lemma 14 implies $v_{i}(B)\leq 4\alpha/3$. Also by Observation 3, $v_{i}(B_{1})\leq 3\alpha/2$ and we have $v_{i}(B_{2})\leq 4\alpha/3$. Therefore
@@
If $|B_{2}|\geq 3$, then by Lemma 14, $v_{i}(B_{2})\leq 4\alpha/3$. Also if $|B_{2}|=1$, $v_{i}(B_{2})<\alpha$ since the instance is $R^{\alpha}_{0}$ irreducible. Therefore, $|B_{2}|=2$. It implies $|B_{1}|=2$ and we have $B_{1}=\{1,n+1\}$ and $B_{2}=\{2,n+2\}$. Let $a$ and $b$ be the agents who received $B_{1}$ and $B_{2}$. If $v_{i}(\{1,n+3\})\geq\alpha$, then the instance is not $T^{\alpha}$-irreducible since $a,b,i$ satisfy its precondition. Therefore, $v_{i}(1)+v_{i}(n+3)<\alpha$.
@@
$<(\alpha/2-\alpha/3)$ ($v_{i}(n+2)\leq v_{i}(n+1)<\alpha/2$ due to Observation 1)
@@
By Lemma 15, the bags $B_{1},\ldots,B_{n-n^{\prime}}$ are allocated during the round robin phase. By Lemma 21, there exists at least $2n/3\geq n^{\prime}$ many bags $B_{i_{1}},\ldots,B_{i_{n^{\prime}}}$ for $i_{1}<i_{2}<\ldots<i_{n^{\prime}}\leq n-n^{\prime}$ that were allocated to other agents during the round robin phase and $v_{i}(B_{i_{\ell}})>\alpha$ for all $\ell\in[n^{\prime}]$. Note that $i_{n^{\prime}}\geq n^{\prime}$. Since $B_{j}$ have been assigned a good in all but maybe the last round, we have $|B_{j}|\geq|B_{i_{n^{\prime}}}|-1$. Let $B_{j}=\{g_{1},\ldots,g_{k}\}$ and $B_{i_{n^{\prime}}}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k^{\prime}}\}$ for $j=g_{1}<\ldots<g_{k}$ and $i_{n^{\prime}}=g^{\prime}_{1}<\ldots<g^{\prime}_{k^{\prime}}$. For all $\ell\in[k^{\prime}-1]$, since $g_{\ell}$ is allocated before $g^{\prime}_{\ell+1}$, $g_{\ell}<g^{\prime}_{\ell+1}$ and hence $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell+1})$. We have
@@
Theorem 5.
Given an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with $n^{*}>5$ agents and $\alpha\leq 4/5$,
$\mathsf{BagFill\mbox{-}RoundRobin}(\mathtt{T-reduce(\mathcal{I},\alpha)},\alpha)$
returns a simple $\alpha$-MMS allocation with at most $\lfloor n^{*}/3\rfloor$ many distinct copies.
@@
Let $n$ be the number of agents after applying $\mathtt{T-reduce(\mathcal{I},\alpha)}$. Since $T^{\alpha}$ and $R^{\alpha}_{k}$ are valid reductions for all $k<|M|/n^{*}$, all agents who receive a bag during the $\mathtt{T-reduce(\mathcal{I},\alpha)}$ receive at least $\alpha$ fraction of their MMS value and the MMS value of the remaining agents do not decrease after this process. Without loss of generality, let the MMS value of the remaining agents be $1$ at this point of the algorithm. All the agents who receive a bag during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}$ get at least $\alpha$ fraction of their value. Let agent $i$ be an agent who receives bag $B_{j}$ after the round robin phase. Let $g$ be the good that was duplicated and added to $B_{j}$ right before allocating it to agent $i$. Since $g\leq n^{\prime}$, $v_{i}(g)\geq v_{i}(n^{\prime})$. By Lemma 22, $v_{i}(B_{j}\cup\{g\})\geq\alpha$. Therefore, the final allocation is $\alpha$-MMS.
We now give an upper bound on the total number of copies and prove that the generated MMS allocation is simple, namely that there are $t\leq\lfloor n^{*}/3\rfloor$ distinct copies, and that the $2t$ instances, composed of these copies and their corresponding original goods, are allocated to $2t$ distinct agents. Let $k=n^{*}-n$ be the number of agents who receive a bag in $\mathtt{T-reduce(\mathcal{I},\alpha)}$. Only when applying $T^{\alpha}$, new duplications are introduced and one duplication per three agents that are removed. So during $\mathtt{T-reduce(\mathcal{I},\alpha)}$, at most $\lfloor k/3\rfloor$ distinct copies are used. Each instance of these good is allocated to a different agent. Furthermore, all the instances of the goods that are allocated during $T^{\alpha}$ along with their owners, are removed from the instance. Therefore, these goods cannot be copied more than once and their owners cannot get any other goods later on. In the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}$, we duplicate items $1,2,\ldots,n^{\prime}$ once and allocate the original and the duplications to $2n^{\prime}$ different agents. By Corollary 6, $n^{\prime}\leq\lfloor n/3\rfloor$. Therefore, in total at most $\lfloor n^{*}/3\rfloor$ distinct copies are used and the allocation is simple. ∎
Appendix A Full MMS by Removing $n-2$ Bads for Chore Allocation
Consider the case where for an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, $i\in N$ and $j\in M$, $v_{ij}$ is the cost of performing chore $j$. We assume agents are additive, that is the cost of bundle $S\subseteq M$ for agent $i\in N$ equals $v_{i}(S)=\sum_{j\in S}v_{ij}.$ Agents want to minimize their cost.
Recall that we denote the collection of all allocations of the set $M$ among $n$ agents by $\mathcal{A}_{n}(M)$. The maximin share [^4] Although the maximin share is defined with a min-max type formula, if one considers chores to be goods with negative value, then a max-min formulation is equivalent. (MMS) of agent $i$ is the upper bound on the cost agent $i$ may exert by splitting the set $M$ into $n$ bundles, getting the worst out of them. Formally:
Definition 8 (Maximin share (MMS)).
For an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with $n$ agents, the maximin share (MMS) of agent $i$ with cost function $v_{i}$, denoted by $\mu_{i}^{n}(M)$, is given by
$\mu_{i}^{n}(M)=\min_{A\in\mathcal{A}_{n}(M)}\max_{S\in A}v_{i}(S).$
For every agent $i$, we normalize $\mu_{i}^{n}(M)=1$, thus $v_{i}(M)\leq n$ for every $i\in N$. We next present Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill\mbox{-}for\mbox{-}Chores}$, which operates in an analogous way to Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ from Section 4.1.
Let $P=(P_{1},P_{2},\ldots,P_{n})$ be the MMS partition of agent 1
Consider the bipartite graph $G=(V=N\times P,E)$, where $E=\{(i,P_{j})\ :\ v_{i}(P_{j})\leq 1\}$
if there is a perfect matching between $N$ and $P$ in $G$ then
       Allocate each $i\in N$ its match in the perfect matching
end if
else
       Let $P^{\prime}\subset P$ be the minimal set of bundles violating Hall’s theorem ($|P^{\prime}|>1$ since $1$ is matched to every node in $P$)
       Find a matching between bundles in $\hat{P}$ and agents in $N$
       foreach $i\in N$ matched to some $P_{j}\in\hat{P}$ do
             $A_{i}\leftarrow P_{j}$
       end foreach
       Allocate agents in $\tilde{N}$ using procedure $\mathsf{BagFill\mbox{-}and\mbox{-}Remove}((\tilde{N},\tilde{M},\{v_{i}\}_{i\in\tilde{N}}))$
end if
Input: $n$ additive cost functions $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at least $m-n+2$ allocated items.
Algorithm 7 $\mathsf{Match\mbox{-}n\mbox{-}Fill\mbox{-}for\mbox{-}Chores}(\mathcal{I}=(N,M,\{v_{i}\}_{i\in N}))$:
Input: $n$ additive cost functions $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at least $m-n+2$ allocated items.
$t\leftarrow 0$, $B_{t}\leftarrow\emptyset$
$N_{0}\leftarrow N$, $M_{0}\leftarrow M$
for items $j\leftarrow 1$ to $m$ do
       if $\exists i\in N_{t}$ such that $v_{i}(B_{t}\cup\{j\})<1$ then
             $B_{t}\leftarrow B_{t}\cup\{j\}$
       end if
      else
             Let $i_{t}\in\operatorname*{arg\,min}_{i\in N_{t}}v_{i}(B_{t})$
             $A_{i_{t}}\leftarrow B_{t}$
             $t\leftarrow t+1$, $B_{t}\leftarrow\emptyset$
             $N_{t}\leftarrow N_{t-1}\setminus\{i_{t-1}\}$, $M_{t}\leftarrow M_{t-1}\setminus(B_{t-1}\cup\{j\})$
       end if
end for
Let $i_{t}$ be the last remaining agent in $N_{t}$. Set $A_{i_{t}}\leftarrow B_{t}$.
Input: $n$ additive cost functions $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at least $|M|-|N|+1$ allocated items.
Algorithm 8 $\mathsf{BagFill\mbox{-}and\mbox{-}Remove}(\mathcal{I}=(N,M,\{v_{i}\}_{i\in N}))$:
Input: $n$ additive cost functions $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at least $|M|-|N|+1$ allocated items.
We first show that $\mathsf{BagFill\mbox{-}and\mbox{-}Remove}$ never gives an agent more than their MMS value, and throws away at most $|N|-1$ items.
Lemma 23.
Assuming that for every $i\in N$, $v_{i}(M)\leq|N|$, $\mathsf{BagFill\mbox{-}and\mbox{-}Remove}$ outputs and allocation such that $v_{i}(A_{i})\leq 1$ for every $i\in N$, and allocates at least $|M|-|N|+1$ items.
Proof.
Every time an agent is allocated, except the last agent, we throw away the current item considered. Thus, we allocate all items but $|N|-1$. Since every time we allocate, we remove a set of items that is worth at least 1 for all agents from $M_{t}$, we maintain the invariant that $v_{i}(M_{t})\leq|N|-t$. Thus, when we reach the last agent, they get a bundle of cost at most 1. ∎
We now show that at most $n-2$ items should be discarded.
Lemma 24.
For any fair division instance of chores $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with additive costs, Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill\mbox{-}for\mbox{-}Chores}(\mathcal{I})$ outputs a full MMS allocation and allocates at least $|M|-|N|+2$ items.
Proof.
As in the proof of Lemma 4, if all agents are allocated by a perfect matching in $G$, and then we do not need to throw away any items. If we match a subset of the agents by bundles in $G$, then the remaining agents do not have any edges to the allocated bundles, which means that if $k$ agents were allocated, the rest of the agents have a cost of at most $|N|-k$ for the remaining bundles. Thus, we can apply Lemma 23 and get that at most $|N|-k-1$ items are thrown away. Since $k\geq 1$, we get that the number of allocated items is at least $|M|-|N|+2$. ∎
Appendix B Missing Proofs of Section 3
Proof of Lemma 2.
@@ Appendix B Missing Proofs of Section 3
For any agent $i$, let $P^{i}=(P^{i}_{1},\dots,P^{i}_{n})$ be his MMS partition. Consider the MMS allocations $A=(A_{1},\dots,A_{n})$ where $A_{i}\sim U(\{P^{i}_{1},\dots,P^{i}_{n}\})$. Namely, $i$’s bundle, $A_{i}$, is one of his MMS bundles chosen uniformly at random, independent of other choices. We use the balls-and-bins analogy [38], and consider the process of constructing $A$ as throwing $n$ balls (the agents) into $n$ bins (the index in an MMS partition). We are interested in the number of balls per bin, so we ignore the fact that the balls are distinct.
@@
Fix some good $g\in M$ and assume without loss that $g\in P^{i}_{1}$ for any $i\in N$. The event in which $g$ is copied at least $k$ times for the allocation $A=(A_{1},\dots,A_{n})$, is exactly the event in which the bin corresponding to index $1$ has at least $k$ balls. A simple analysis shows that,
@@
$\Pr[g\text{ was copied at least }k\text{ times}]\leq\binom{n}{k}\left(\frac{1}{n}\right)^{k}\leq\left(\frac{e}{k}\right)^{k},$
@@
Pr[there exists a good with more than $\Pr\left[\text{there exists a good with more than }\frac{3\ln m}{\ln\ln m}\text{ copies}\right]<m\cdot\frac{1}{m^{2}}=\frac{1}{m},$
@@
Appendix C Missing Proofs of Section 4
Proof of Lemma 3.
Every time an agent is allocated, except the last agent, we duplicate the last item that was put in his bag and put the duplicate as the first item in the next bag. Since every item’s value is less than $1$ for all agents, every allocated bundle contains at least two items. Therefore, we duplicate a different item each time, and in total, $|N|-1$ items. We now show that upon termination, every agent gets at least their MMS value.

### v3

@@ 4.3 Combining Round Robin and Bag-Filling
By Lemma 4.19, the bags $B_{1},\ldots,B_{n-n^{\prime}}$ are allocated during the round robin phase. By Lemma 4.22, there exists at least $n^{\prime}$ many bags $B_{i_{1}},\ldots,B_{i_{n^{\prime}}}$ for $i_{1}<i_{2}<\ldots<i_{n^{\prime}}\leq n-n^{\prime}$ that were allocated to other agents during the round robin phase and $v_{i}(B_{i_{\ell}})>1$ for all $\ell\in[n^{\prime}]$. Note that $i_{n^{\prime}}\geq n^{\prime}$. Since $B_{j}$ have been assigned a good in all but maybe the last round, we have $|B_{j}|\geq|B_{i_{n^{\prime}}}|-1$. Let $B_{j}=\{g_{1},\ldots,g_{k}\}$ and $B_{i_{n^{\prime}}}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k^{\prime}}\}$ for $j=g_{1}<\ldots<g_{k}$ and $i_{n^{\prime}}=g^{\prime}_{1}<\ldots<g^{\prime}_{k^{\prime}}$. For all $\ell\in[k^{\prime}-1]$, since $g_{\ell}$ is allocated before $^{\prime}_{\ell+1}, g<g$$_{\ell}<g^{\prime}_{\ell+1}$ and hence $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell+1})$. We have
@@
Proof of Theorem 4.2.
We prove that for ordered instances $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$,
@@
returns a simple $\alpha$-MMS allocation with at most $\lfloor n^{*}/2\rfloor$ many distinct copies. Then, from Proposition 4.1, the theorem follows.
Let $n$ be the number of agents after applying $\mathtt{S-reduce(\mathcal{I},\alpha)}$. Since $S^{\alpha}$ and $R^{\alpha}_{k}$ are valid reductions for all $k<|M|/n^{*}$, all agents who receive a bag during the $\mathtt{S-reduce(\mathcal{I},\alpha)}$ receive at least $\alpha$ fraction of their MMS value and the MMS value of the remaining agents do not decrease after this process. Without loss of generality, let the MMS value of the remaining agents be $1$ at this point of the algorithm. All the agents who receive a bag during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}$ get at least $\alpha$. Let agent $i$ be an agent who receives bag $B_{j}$ after the round robin phase. Let $g$ be the good that was duplicated and added to $B_{j}$ right before allocating it to agent $i$. $g\leq n^{\prime}$ and thus $v_{i}(g)\geq v_{i}(n^{\prime})$. By Lemma 4.24, $v_{i}(B_{j}\cup\{g\})\geq\alpha$. Therefore, the final allocation is $\alpha$-MMS.
We now prove that the generated MMS allocation is simple, namely that there are $t\leq\lfloor n^{*}/2\rfloor$ distinct copies, and that the $2t$ instances, composed of these copies and their corresponding original goods, are allocated to $2t$ distinct agents. Let $k=n^{*}-n$ be the number of agents who receive a bag in $\mathtt{S-reduce(\mathcal{I},\alpha)}$. Only when applying $S^{\alpha}$, new duplications are introduced and one duplication per two agents that are removed. So during $\mathtt{S-reduce(\mathcal{I},\alpha)}$, at most $\lfloor k/2\rfloor$ distinct copies are used. Each instance of these good is allocated to a different agent. Furthermore, all the instances of the goods that are allocated during $S^{\alpha}$ along with their owners, are removed from the instance. Therefore, these goods cannot be copied more than once and their owners cannot get any other goods later on. In the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}$, we duplicate items $1,2,\ldots,n^{\prime}$ once and allocate the original and the duplications to $2n^{\prime}$ different agents. By Corollary 4.29, $n^{\prime}\leq\lfloor n/2\rfloor$. Therefore, in total at most $\lfloor n^{*}/2\rfloor$ distinct copies are used and the allocation is simple. ∎
[…]
In this section, we present an algorithm which, given an ordered instance, outputs a simple $\alpha$-MMS allocation with at most $\lfloor n/3\rfloor$ distinct copies for any $\alpha\leq 4/5$. Then, using Proposition 4.1, we prove the following theorem in Appendix C.
See 4.3
@@
Input: An ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and approximation factor $\alpha$
@@
Input: An ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and approximation factor $\alpha$
@@
The lemma follows from Lemma 4.17 and Corollary 4.10. ∎
For the rest of this (sub)-section, let $\alpha\leq 4/5$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible and $T^{\alpha}$-irreducible instance, and $\mu_{i}(\mathcal{I})=1$. Let $n$ be the number of agents right before the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, $\mathcal{B}=\{B_{1},\ldots,B_{a-1}\}$ be the set of bags allocated during this phase, and $n^{\prime}=n-a+1$ be the number of remaining agents after this phase.
Lemma 4.26.
@@
Lemma 4.28.
Let $i$ be a remaining agent after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ (if any). Then $|\{B\in\mathcal{B}\mid v_{i}(B)>\alpha\}|\geq 2n/3$ or $n\leq 5$.
@@
Assume $n\geq 6$. Let $\mathcal{B}^{+}=\{B\in\mathcal{B}\mid v_{i}(B)>\alpha\}$. By definition, for all $B\in\mathcal{B}\setminus\mathcal{B}^{+}$, $v_{i}(B)\leq\alpha$. Let $\bar{\mathcal{B}}$ be the set of $n-|\mathcal{B}|=n^{\prime}$ many remaining bags after the round robin phase. For all $B\in\bar{\mathcal{B}}$, $v_{i}(B)<\alpha$. Otherwise, $B$ should have been allocated to $i$ (or some other agent who values it at least $\alpha$ fraction of their MMS value). We consider two cases:
@@
For all $B\in\mathcal{B}\setminus\{B_{1},B_{2}\}$, Lemma 4.26 implies $|B|\geq 3$ and Lemma 4.18 implies $v_{i}(B)\leq 4\alpha/3$. Also by Observation 4.27, $v_{i}(B_{1})\leq 3\alpha/2$ and we have $v_{i}(B_{2})\leq 4\alpha/3$. Therefore
@@
If $|B_{2}|\geq 3$, then by Lemma 4.18, $v_{i}(B_{2})\leq 4\alpha/3$. Also if $|B_{2}|=1$, $v_{i}(B_{2})<\alpha$ since the instance is $R^{\alpha}_{0}$ irreducible. Therefore, $|B_{2}|=2$. It implies $|B_{1}|=2$ and we have $B_{1}=\{1,n+1\}$ and $B_{2}=\{2,n+2\}$. Let $a$ and $b$ be the agents who received $B_{1}$ and $B_{2}$. If $v_{i}(\{1,n+3\})\geq\alpha$, then the instance is not $T^{\alpha}$-irreducible since $a,b,i$ satisfy its precondition. Therefore, $v_{i}(1)+v_{i}(n+3)<\alpha$.
@@
$<(\alpha/2-\alpha/3)$ ($v_{i}(n+2)\leq v_{i}(n+1)<\alpha/2$ due to Observation 4.12)
@@
By Lemma 4.19, the bags $B_{1},\ldots,B_{n-n^{\prime}}$ are allocated during the round robin phase. By Lemma 4.28, there exists at least $2n/3\geq n^{\prime}$ many bags $B_{i_{1}},\ldots,B_{i_{n^{\prime}}}$ for $i_{1}<i_{2}<\ldots<i_{n^{\prime}}\leq n-n^{\prime}$ that were allocated to other agents during the round robin phase and $v_{i}(B_{i_{\ell}})>\alpha$ for all $\ell\in[n^{\prime}]$. Note that $i_{n^{\prime}}\geq n^{\prime}$. Since $B_{j}$ have been assigned a good in all but maybe the last round, we have $|B_{j}|\geq|B_{i_{n^{\prime}}}|-1$. Let $B_{j}=\{g_{1},\ldots,g_{k}\}$ and $B_{i_{n^{\prime}}}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k^{\prime}}\}$ for $j=g_{1}<\ldots<g_{k}$ and $i_{n^{\prime}}=g^{\prime}_{1}<\ldots<g^{\prime}_{k^{\prime}}$. For all $\ell\in[k^{\prime}-1]$, since $g_{\ell}$ is allocated before $^{\prime}_{\ell+1}, g<g$$_{\ell}<g^{\prime}_{\ell+1}$ and hence $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell+1})$. We have
@@
5 Discussion
Herein, we show that exact MMS fairness can be guaranteed in settings that permit post facto adjustments to the supply of indivisible items. In particular, we obtain tight bounds for XOS valuations and beyond, and, in the context of $\alpha$-MMS, improve upon the state of the art using few copies.
The framework of resource augmentation provides a meaningful route for bypassing barriers in discrete fair division. In particular, it would be interesting to develop truthful mechanisms for MMS that leverage bounded supply adjustment. Another relevant direction for future work is to consider the framework for achieving envy-freeness up to any good (EFX). Such an approach would complement prior works on EFX with charity (see, e.g., [chaudhury2021little]), which obtain existential guarantees for EFX with disposal, rather than duplication, of goods.
One may consider several metrics for measuring the augmentation in an allocation with copies. In this work we consider the number of copies made of each good and the total numbers of copies made ($k$ and $t$, respectively, in $(t,k)$-allocation). Another possible measure is the total number of allocated goods (including copies), required to achieve MMS. This notion is relevant for settings where unallocated goods are not manufactured (e.g., course allocation), or are refunded (e.g., medical supplies). Budish [budish2011combinatorial] considered the $\ell_{2}$ norm of the $m$-dimensional vector whose components correspond to the number of copies made of the $m$ goods, respectively. Clearly one can consider many other measures which suggest a host of new problems.
Appendix A Proof of 4
Towards contradiction, assume there exists $f:\mathbb{N}\rightarrow\mathbb{N}$ such that 11-out-of-$f(n)$ MMS allocations exist for all instances with $n$ agents. Let $n^{\prime}=f(n)$ and consider the $n^{\prime}$-dimensional cube and its corresponding XOS valuations as in Theorem 3.1. We call this instance $\mathcal{I}=\langle[n^{\prime}],M,\{v_{i}\}_{i\in[n^{\prime}]}\rangle$. Let $N\subset[n^{\prime}]$ be the subset of the first $n$ agents in this example; $N=[n]$. Let $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$. For all $i\in N$, $\mu^{f(n)}_{i}(M)=\mu_{i}(\mathcal{I}^{\prime})$. Therefore, there exists an allocation $A=(A_{1},\ldots,A_{n})$ such that for all $i\in[n]$
$v_{i}(A_{i})\geq\mu^{f(n)}_{i}(M)=\mu_{i}(\mathcal{I}^{\prime}).$
Hence, $P^{1}_{i}\subseteq A_{1}$ and $P^{2}_{j}\subseteq A_{2}$ for some $i,j\in[n^{\prime}]$, and $A_{1}\cap A_{2}=\emptyset$. However, $P^{1}_{i}\cap P^{2}_{j}=\{g_{d_{1},\dots,d_{n}}\mid\;d_{1}=i,d_{2}=j\}\neq\emptyset$ which is a contradiction.
Appendix B Missing Proofs of Section 3
See 3.5
@@ Appendix B Missing Proofs of Section 3
For any agent $i$, let $P^{i}=(P^{i}_{1},\dots,P^{i}_{n})$ be his MMS partition. Consider the MMS allocation $A=(A_{1},\dots,A_{n})$ where bundle $A_{i}\sim\mathrm{Unif}(\{P^{i}_{1},\dots,P^{i}_{n}\})$, independently for each agent $i$. That is, $i$’s bundle, $A_{i}$, is one of his MMS subsets chosen uniformly at random, independent of other choices. We use the balls-and-bins analogy [raab1998balls], and consider the process of constructing $A$ as throwing $n$ balls (the agents) into $n$ bins (the index in an MMS partition). We are interested in the number of balls per bin, so we ignore the fact that the balls are distinct.
@@
Fix some good $g\in M$ and assume without loss that $g\in P^{i}_{1}$ for any $i\in N$. The event in which $g$ is copied at least k1k-1 times for the allocation $A=(A_{1},\dots,A_{n})$, is exactly the event in which the bin corresponding to index $1$ has at least $k$ balls. A simple analysis shows that,
@@
$\Pr[g\text{ was copied at least }k-1\text{ times}]\leq\binom{n}{k}\left(\frac{1}{n}\right)^{k}\leq\left(\frac{e}{k}\right)^{k},$
@@
[…]
@@
See 3.7
Proof.
Every time an agent is allocated, except the last agent, we duplicate the last item that was put in his bag and put the duplicate as the first item in the next bag. Since the value of every item is less than $1$ for all agents, every allocated bundle has at least two items. Thus, we duplicate a different item each time, and in total, |N|1|N|-1 items. We now show that upon termination, every agent gets at least a value of 1.